# Inversion sequences growing on the left

### v2

@@ Abstract
Generating trees growing on the left for pattern-avoiding inversion sequences
Benjamin Testart [^1] Université de Lorraine, CNRS, Inria, LORIA, F-54000 Nancy, France
Keywords: pattern avoidance, inversion sequences, enumeration, generating trees, generating functions
@@ 1.1 Basic definitions
For all $n\geqslant 0$, let $\mathcal{I}_{n}$ be the set of inversion sequences [MS15, CMSW16] of size $n$, that is, the set of sequences $\sigma=(\sigma_{1},\dots,\sigma_{n})$ such that $\sigma_{i}\in\{0,\dots,i-1\}$ for all $i\in\{1,\dots,n\}$. Let $\mathcal{I}=\coprod_{n\geqslant 0}\mathcal{I}_{n}$ be the set of all inversion sequences.
@@ 1.2 Context and summary of results
Pattern-avoiding inversion sequences were introduced in [MS15] and [CMSW16], by analogy with pattern-avoiding permutations. Afterwards, many more articles studied inversion sequences avoiding patterns of length three, see [MS18, BBGR19, YL20, KMY24, CL24, Pan24, Tes24], among others.
Our work is concerned with two distinct classes of pattern-avoiding inversion sequences: $\mathcal{I}(201,210)$ and $\mathcal{I}(010,102)$. The class $\mathcal{I}(201,210)$ is counted by sequence A212198 in [OEIS24]. It has an algebraic generating function, which was found in both [CL24] and [Pan24], using two different constructions. Chen and Lin [CL24] use a decomposition around the rightmost saturated [^2] In an inversion sequence $\sigma$, the entry $\sigma_{i}$ is said to be “saturated” if $\sigma_{i}=i-1$ (since i1i-1 is the largest value which could appear at position $i$). entry of inversion sequences, with one parameter counting the number of saturated entries. The construction used by Pantone [Pan24] is a generating tree in one integer parameter and two boolean parameters, and it involves some “commitments” to future decisions.
The class $\mathcal{I}(010,102)$ is counted by sequence A374553 in [OEIS24]. This sequence was computed in polynomial time in [Tes24] using a decomposition involving $\{010,102\}$-avoiding words, and it was conjectured to have an algebraic generating function in [Pan24] using the software [Pan21].
In Section 2, we describe a construction for inversion sequences, which we call the “generating tree growing on the left”. This construction first appeared in [CMSW16] to enumerate the classes $\mathcal{I}(101)$ and $\mathcal{I}(110)$. A modified construction based on this one also appears in [LF21] to enumerate inversion sequences avoiding the vincular pattern $\underline{12}0$.
As it is often the case with generating trees, our general approach in this paper is to first look for a simple succession rule describing the tree (up to isomorphism), then turn this rule into a functional equation for a multivariate generating function, then solve this equation with the kernel method [BBDFGG02].
@@
In Section 4, we modify the generating tree growing on the left construction to apply it to inversion sequences avoiding 010 and 102. In this case, we obtain a functional equation involving two catalytic variables, which can be solved by simply applying the kernel method twice. This allows us to compute the generating function of $\mathcal{I}(010,102)$, confirming a conjecture of [Pan24].
@@ 2 Generating trees for inversion sequences growing on the left
A generating tree [Wes95, Wes96] is a rooted, labelled tree such that the label of each node determines its number of children, and their labels. A generating tree can be defined by a succession rule
@@
Generating trees are used to solve enumeration problems by identifying combinatorial objects with the nodes of a tree, so that each object of size $n$ corresponds to a node at level $n$ in the tree. This is useful when many nodes are roots of isomorphic subtrees, as the combinatorial objects corresponding to these nodes may all be given the same label. A succession rule for a generating tree then induces a recurrence relation on the number of objects of each size (which usually involves some parameters corresponding to the labels of the tree), or equivalently a functional equation for a (multivariate) generating function. For a more thorough introduction to generating trees in the context of pattern-avoiding inversion sequences, see [Tes24, Section 2.1].
@@
Observe that this construction is indeed a generating tree (it produces only inversion sequences and, starting from the empty sequence $\varepsilon$, it produces every inversion sequence exactly once). Note that steps 1 and 3 are deterministic, whereas step 2 involves choosing a subset of the zero entries. For any inversion sequence $\sigma\in\mathcal{I}$, and any $Z\subseteq\mathbf{Zeros}(\sigma)$, we denote by $\mathbf{child}(\sigma,Z)=0\cdot(\sigma_{i}+\chi(\sigma_{i}>0)+\chi(i\in Z))_{i\in[1,|\sigma|]}$ the child of $\sigma$ corresponding to the subset of zeros $Z$ to which 1 is added. It should be apparent that $\sigma$ has $2^{|\mathbf{Zeros}(\sigma)|}$ children. For instance, the sequence $\sigma=(0,1,1,3,0,2,5)$ satisfies $\mathbf{Zeros}(\sigma)=\{1,5\}$, and its four children are
@@
The simplest construction of inversion sequences is the generating tree growing sequences on the right, used in e.g. [KMY24] or [Tes24, Section 2]. In comparison, growing sequences on the left is not the most practical way to construct the set of inversion sequences (it is not even obvious that the tree described by $\Omega_{\text{Left}}$ has $n!$ nodes at each level $n$). However, this construction becomes simpler for some classes of pattern-avoiding inversion sequences. For instance, [CMSW16] proves that the generating trees growing on the left for the classes $\mathcal{I}(101)$ and $\mathcal{I}(110)$ are both described by the succession rule
@@ 4.1 Modifying the construction
The article [FPPR03] introduces the notion of “jumping” generating trees, in which each edge has a (bounded) positive integer length. If a node at level $n$ has a child through an edge of length $k$, then this child is at level $n+k$ (in a classical generating tree, all edges have length $1$). The tree in Figure 2 could be seen as a jumping generating tree having edges of length 0 or 1.
@@
  • The letter $\mathtt{L}$ marks sequences whose entries of value 1 are on the left, of the form $(0,\dots,0,1\dots,1)\cdot\mu$.
  • The letter $\mathtt{R}$ marks sequences whose entries of value 1 are on the right, of the form $(0,\dots,0)\cdot\mu\cdot(1,\dots,1)$.
  • The letter $\mathtt{M}$ marks sequences whose entries of value 1 are in the middle, of the form $(0,\dots,0)\cdot\mu\cdot(1,\dots,1)\cdot\nu$.
@@ 5 A broader perspective on our construction
Pattern Succession rule Comment [OEIS24]
@@
011 $\begin{array}[]{rclll}(0)\\ (z)&\leadsto&(z)^{z},\,(z+1)\end{array}$ Isomorphic to the generating tree growing on the right used in [CMSW16].
@@
012 $\begin{array}[]{rclll}(\mathtt{A},0)\\ (\mathtt{A},z)&\leadsto&(\mathtt{A},z+1)\\ &&(\mathtt{B},i)^{2^{z-1-i}}&\text{for}&i\in[0,z-1]\\ (\mathtt{B},s)&\leadsto&(\mathtt{B},s)\\ &&(\mathtt{B},i)^{2^{s-1-i}}&\text{for}&i\in[0,s-1]\end{array}$ The construction of [CMSW16] is much simpler (it can be seen as a generating tree with only two distinct labels).
@@
Fibonacci)
021 $\begin{array}[]{rclll}(0)\\ (p)&\leadsto&(p+1)\\ &&(i)^{2^{p-i}}&\text{for}&i\in[1,p]\end{array}$ See [CMSW16] for two bijections between $\mathcal{I}(021)$ and Schröder paths, and a bijection with some binary trees. See also [LK18] for a bijection with two-colored Dyck paths.
@@
Schröder)
$\begin{array}[]{c}101\\ 110\end{array}$ $\begin{array}[]{rclll}(0)\\ (z)&\leadsto&(z+1)\\ &&(i)^{i}&\text{for}&i\in[1,z]\end{array}$ This construction is used in [CMSW16] and [BBGR19]. A113227
102 $\begin{array}[]{l}\begin{array}[]{rclll}(0,0)\\ (z,z)&\leadsto&(z+1,z+1)\\ &&(i,j)^{\binom{z-i-1}{z-j}}&\hskip 23.0pt\text{for}&i\in[0,z-1],\,j\in[i+1,z]\end{array}\\ \text{if}\;s\neq z:\\ \begin{array}[]{rclll}(s,z)&\leadsto&(s,j)&\text{for}&j\in[s+1,z+1]\\ &&(i,j)^{\sum_{k=s-j}^{z-j}\binom{s-i-1}{k}}&\text{for}&i\in[0,s-1],\,j\in[i+1,z]\end{array}\end{array}$ A different construction with two parameters is used in [MS15] to compute the generating function of $\mathcal{I}(102)$. A200753
$\begin{array}[]{c}201\\ 210\end{array}$ $\begin{array}[]{rclll}(0,0)\\ (p,z)&\leadsto&(p+1,j)&\text{for}&j\in[p+1,z+1]\\ &&(i,j)^{\sum_{k=p-j}^{z-j}\binom{p-i}{k}}&\text{for}&i\in[1,p],\,j\in[i,z]\end{array}$ Two other constructions for this class, involving two parameters each, can be found in [CMSW16] and [MS15]. A263777
The number of distinct labels appearing at level $n$ of the tree described by the rule $\Omega_{120}$ is exponential in $n$. In this case, the approach of [MS15] is much better since it can be used to compute the enumeration sequence of $\mathcal{I}(120)$ in polynomial time (this is sequence A263778 in the OEIS). We present some more serviceable succession rules for 120-avoiding inversion sequences ($\Omega^{\prime}_{120}$ and $\Omega^{\prime\prime}_{120}$) towards the end of this section.
Lastly, we show that even for the classes $\mathcal{I}(000)$, $\mathcal{I}(010)$, and $\mathcal{I}(100)$, it is possible to modify the generating tree growing on the left to obtain some succession rules, in which some of the labels correspond to objects that are not counted (such objects are referred to as “phantom objects” in [Tes24]). The OEIS entries related to these three classes are A000111, A263780, and A263779.
@@
Inversion sequences avoiding a constant pattern of any length have been studied before, and their exponential generating function is given in [HL22]. We have not tried to compute generating functions from the rule $\Omega_{0^{k}}$.
@@
By applying this idea to the pattern 010, we can obtain a generating tree having similar labels to those of the rule $\Omega_{120}$ (to be precise, each inversion sequence is labelled by the number of zeros in its prefix, and the multiset of lengths of the other factors of zeros). The number of distinct labels at level $n$ of this tree is exponential in $\sqrt{n}$, so this does not provide an efficient way of computing $|\mathcal{I}_{n}(010)|$. Nevertheless, it is possible to improve this construction to obtain a tree which only has a quadratic number of distinct labels at level $n$, by using the idea of commitments from [Pan24]. That tree is described by the succession rule below, where the terms in brackets are unsigned Stirling numbers of the first kind, and only nodes labelled $(p,0)$ must be counted to enumerate $\mathcal{I}(010)$.
@@
With this rule, the enumeration sequence of $\mathcal{I}(010)$ can be computed with a similar efficiency to the method previously used in [Tes24].
@@
The rule $\Omega^{\prime}_{120}$ has a nicer form than $\Omega_{010}$, making it possible to skip the commitments. This yields the rule $\Omega^{\prime\prime}_{120}$ below, which only has one parameter, although it requires allowing jumps of arbitrarily long length. As a consequence, some nodes of the corresponding “tree” have an infinite number of children. This construction of $\mathcal{I}(120)$ is actually quite similar [^6] The construction of $\mathcal{I}(120)$ from [MS15] works by splitting an inversion sequence into factors starting at each left-to-right maximum, and building each factor from left to right. Here, we essentially build these same factors from right to left. to the one employed in [MS15].
@@
In conclusion, up to allowing some small modifications, a succession rule involving at most two integer parameters can be found to describe a kind of generating tree growing on the left for every class of inversion sequences avoiding a pattern of length 3. This shows that the construction is quite robust, and it could be further explored in a larger setting; for instance, to study longer patterns, sets of patterns, or different definitions of patterns. We also expect this construction to work well with pattern-avoiding restricted growth functions [CDD+18].
Acknowledgements

### v3

# Generating trees growing on the left for pattern-avoiding inversion sequences

@@ Abstract
keywords:
pattern avoidance, inversion sequences, enumeration, generating trees, generating functions
@@ 1.1 Basic definitions
For all $n\geqslant 0$, let $\mathcal{I}_{n}$ be the set of inversion sequences [10, 4] of size $n$, that is, the set of sequences $\sigma=(\sigma_{1},\dots,\sigma_{n})$ such that $\sigma_{i}\in\{0,\dots,i-1\}$ for all $i\in\{1,\dots,n\}$. Let $\mathcal{I}=\coprod_{n\geqslant 0}\mathcal{I}_{n}$ be the set of all inversion sequences.
@@ 1.2 Context and summary of results
Pattern-avoiding inversion sequences were introduced independently in [10] and [4], by analogy with pattern-avoiding permutations. Afterwards, many more articles studied inversion sequences avoiding patterns of length three, see [11, 1, 17, 7, 3, 13, 14], among others.
Our work is concerned with two distinct classes of pattern-avoiding inversion sequences: $\mathcal{I}(201,210)$ and $\mathcal{I}(010,102)$. The class $\mathcal{I}(201,210)$ is counted by sequence A212198 in [OEIS]. It has an algebraic generating function, which was found in both [3] and [13], using two different constructions. Chen and Lin [3] use a decomposition around the rightmost saturated [^1] In an inversion sequence $\sigma$, the entry $\sigma_{i}$ is said to be “saturated” if $\sigma_{i}=i-1$ (since i1i-1 is the largest value which could appear at position $i$). entry of inversion sequences, with one parameter counting the number of saturated entries. The construction used by Pantone [13] is a generating tree in one integer parameter and two boolean parameters, and it involves some “commitments” to future decisions.
The class $\mathcal{I}(010,102)$ is counted by sequence A374553 in [OEIS]. This sequence was computed in polynomial time in [14] using a decomposition involving $\{010,102\}$-avoiding words, and it was conjectured to have an algebraic generating function in [13] using the software [12].
In Section 2, we describe a construction for inversion sequences, which we call the “generating tree growing on the left”. This construction first appeared in [4] to enumerate the classes $\mathcal{I}(101)$ and $\mathcal{I}(110)$. A modified construction based on this one also appears in [8] to enumerate inversion sequences avoiding the vincular pattern $\underline{12}0$.
As it is often the case with generating trees, our general approach in this paper is to first look for a simple succession rule describing the tree (up to isomorphism), then turn this rule into a functional equation for a multivariate generating function, then solve this equation with the kernel method [BBD+].
@@
In Section 4, we modify the generating tree growing on the left construction to apply it to inversion sequences avoiding 010 and 102. In this case, we obtain a functional equation involving two catalytic variables, which can be solved by simply applying the kernel method twice. This allows us to compute the generating function of $\mathcal{I}(010,102)$, confirming a conjecture of [13].
@@ 2 Generating trees for inversion sequences growing on the left
A generating tree [16, 15] is a rooted, labelled tree such that the label of each node determines its number of children, and their labels. A generating tree can be defined by a succession rule
@@
Generating trees are used to solve enumeration problems by identifying combinatorial objects with the nodes of a tree, so that each object of size $n$ corresponds to a node at level $n$ in the tree. This is useful when many nodes are roots of isomorphic subtrees, as the combinatorial objects corresponding to these nodes may all be given the same label. A succession rule for a generating tree then induces a recurrence relation on the number of objects of each size (which usually involves some parameters corresponding to the labels of the tree), or equivalently a functional equation for a (multivariate) generating function. For a more thorough introduction to generating trees in the context of pattern-avoiding inversion sequences, see [14, Section 2.1].
@@
Observe that this construction is indeed a generating tree (it produces only inversion sequences and, starting from the empty sequence $\varepsilon$, it produces every inversion sequence exactly once). Note that steps 1 and 3 are deterministic, whereas step 2 involves choosing a subset of the zero entries. For any inversion sequence $\sigma\in\mathcal{I}$, and any $Z\subseteq\mathbf{Zeros}(\sigma)$, we denote
$\mathbf{child}(\sigma,Z)=0\cdot(\sigma_{i}+\chi(\sigma_{i}>0)+\chi(i\in Z))_{i\in[1,|\sigma|]}$
the child of $\sigma$ corresponding to the subset of zeros $Z$ to which 1 is added. It should be apparent that $\sigma$ has $2^{|\mathbf{Zeros}(\sigma)|}$ children. For instance, the sequence $\sigma=(0,1,1,3,0,2,5)$ satisfies $\mathbf{Zeros}(\sigma)=\{1,5\}$, and its four children are
@@
The simplest construction of inversion sequences is the generating tree growing sequences on the right, used in e.g. [7] or [14, Section 2]. In comparison, growing sequences on the left is not the most practical way to construct the set of inversion sequences (it is not even obvious that the tree described by $\Omega_{\text{Left}}$ has $n!$ nodes at each level $n$). However, this construction becomes simpler for some classes of pattern-avoiding inversion sequences. For instance, [4] proves that the generating trees growing on the left for the classes $\mathcal{I}(101)$ and $\mathcal{I}(110)$ are both described by the succession rule
@@ 4.1 Modifying the construction
The article [5] introduces the notion of “jumping” generating trees, in which each edge has a (bounded) positive integer length. If a node at level $n$ has a child through an edge of length $k$, then this child is at level $n+k$ (in a classical generating tree, all edges have length $1$). The tree in Figure 2 could be seen as a jumping generating tree having edges of length 0 or 1.
@@
  • The letter $\mathtt{L}$ marks sequences having 1-entries on the left, of the form $(0,\dots,0,1\dots,1)\cdot\mu$.
  • The letter $\mathtt{R}$ marks sequences having 1-entries on the right, of the form $(0,\dots,0)\cdot\mu\cdot(1,\dots,1)$.
  • The letter $\mathtt{M}$ marks sequences having 1-entries in the middle, of the form $(0,\dots,0)\cdot\mu\cdot(1,\dots,1)\cdot\nu$.
@@ 5 A broader perspective on our construction
Pattern Succession rule Comment OEIS [OEIS]
@@
011 $\begin{array}[]{rclll}(0)\\ (z)&\leadsto&(z)^{z},\,(z+1)\end{array}$ Isomorphic to the generating tree growing on the right used in [4].
@@
(Bell
numbers)
012 $\begin{array}[]{rclll}(\mathtt{A},0)\\ (\mathtt{A},z)&\leadsto&(\mathtt{A},z+1)\\ &&(\mathtt{B},i)^{2^{z-1-i}}&\text{for}&i\in[0,z-1]\\ (\mathtt{B},s)&\leadsto&(\mathtt{B},s)\\ &&(\mathtt{B},i)^{2^{s-1-i}}&\text{for}&i\in[0,s-1]\end{array}$ The construction of [4] is much simpler (it can be seen as a generating tree with only two distinct labels).
@@
Fibonacci
numbers)
021 $\begin{array}[]{rclll}(0)\\ (p)&\leadsto&(p+1)\\ &&(i)^{2^{p-i}}&\text{for}&i\in[1,p]\end{array}$ See [4] for two bijections between $\mathcal{I}(021)$ and Schröder paths, and a bijection with some binary trees. See also [9] for a bijection with two-colored Dyck paths.
@@
Schröder
numbers)
$\begin{array}[]{c}101\\ 110\end{array}$ $\begin{array}[]{rclll}(0)\\ (z)&\leadsto&(z+1)\\ &&(i)^{i}&\text{for}&i\in[1,z]\end{array}$ This construction is used in [4] and [1]. A113227
102 $\begin{array}[]{l}\begin{array}[]{rclll}(0,0)\\ (z,z)&\leadsto&(z+1,z+1)\\ &&(i,j)^{\binom{z-i-1}{z-j}}&\hskip 23.5pt\text{for}&i\in[0,z-1],\,j\in[i+1,z]\end{array}\\ \text{if}\;s\neq z:\\ \begin{array}[]{rclll}(s,z)&\leadsto&(s,j)&\text{for}&j\in[s+1,z+1]\\ &&(i,j)^{\sum_{k=s-j}^{z-j}\binom{s-i-1}{k}}&\text{for}&i\in[0,s-1],\,j\in[i+1,z]\end{array}\end{array}$ A different construction with two parameters is used in [10] to compute the generating function of $\mathcal{I}(102)$. A200753
$\begin{array}[]{c}201\\ 210\end{array}$ $\begin{array}[]{rclll}(0,0)\\ (p,z)&\leadsto&(p+1,j)&\text{for}&j\in[p+1,z+1]\\ &&(i,j)^{\sum_{k=p-j}^{z-j}\binom{p-i}{k}}&\text{for}&i\in[1,p],\,j\in[i,z]\end{array}$ Two other constructions for this class, involving two parameters each, can be found in [4] and [10]. A263777
The number of distinct labels appearing at level $n$ of the tree described by the rule $\Omega_{120}$ is exponential in $n$. In this case, the approach of [10] is much better since it can be used to compute the enumeration sequence of $\mathcal{I}(120)$ in polynomial time (this is sequence A263778 in [OEIS]). We present some more serviceable succession rules for 120-avoiding inversion sequences ($\Omega^{\prime}_{120}$ and $\Omega^{\prime\prime}_{120}$) towards the end of this section.
We now show that even for the classes $\mathcal{I}(000)$, $\mathcal{I}(010)$, and $\mathcal{I}(100)$, it is possible to modify the generating tree growing on the left to obtain some succession rules, in which some of the labels correspond to objects that are not counted (such objects are referred to as “phantom objects” in [14]). The enumeration sequences of these three classes can be found in [OEIS] as sequences A000111, A263780, and A263779.
@@
Inversion sequences avoiding a constant pattern of any length have been studied before, and their exponential generating function is given in [6]. We have not tried to compute generating functions from the rule $\Omega_{0^{k}}$.
@@
By applying this idea to the pattern 010, we can obtain a generating tree having similar labels to those of the rule $\Omega_{120}$ (to be precise, each inversion sequence is labelled by the number of zeros in its prefix, and the multiset of lengths of the other factors of zeros). The number of distinct labels at level $n$ of this tree is exponential in $\sqrt{n}$, so this does not provide an efficient way of computing $|\mathcal{I}_{n}(010)|$. Nevertheless, it is possible to improve this construction to obtain a tree which only has a quadratic number of distinct labels at level $n$, by using the idea of commitments from [13]. That tree is described by the succession rule below, where the terms in brackets are unsigned Stirling numbers of the first kind, and only nodes labelled $(p,0)$ must be counted to enumerate $\mathcal{I}(010)$.
@@
With this rule, the enumeration sequence of $\mathcal{I}(010)$ can be computed with a similar efficiency to the method previously used in [14].
@@
The rule $\Omega^{\prime}_{120}$ has a nicer form than $\Omega_{010}$, making it possible to skip the commitments. This yields the rule $\Omega^{\prime\prime}_{120}$ below, which only has one parameter, although it requires allowing arbitrarily long jumps. As a consequence, some nodes of the corresponding “tree” have an infinite number of children. This construction of $\mathcal{I}(120)$ is actually quite similar [^5] The construction of $\mathcal{I}(120)$ from [10] works by splitting an inversion sequence into factors starting at each left-to-right maximum, and building each factor from left to right. Here, we essentially build these same factors from right to left. to the one employed in [10].
@@
In conclusion, up to allowing some small modifications, a succession rule involving at most two integer parameters can be found to describe a kind of generating tree growing on the left for every class of inversion sequences avoiding a pattern of length 3. This shows that the construction is quite robust, and it could be further explored in a larger setting; for instance, to study longer patterns, sets of patterns, or different definitions of patterns. We believe this construction could also be applied to restricted growth functions [CDD+], since they form a subtree of the generating tree growing on the left for inversion sequences: the parent of a (nonempty) restricted growth function is always a restricted growth function. However, this does not hold for ascent sequences [2], another interesting subset of inversion sequences.
Acknowledgements.